Use the sample data and confidence level given below to complete parts​ (a) through​ (d).

A research institute poll asked respondents if they felt vulnerable to identity theft. In the​ poll, n=1025 and x=520 who said​ "yes." Use a 90% confidence level.

a) Find the best point estimate of the population proportion p.

=0.507

b) Identify the value of the margin of error E.

E=
Answer #1

Solution:- Given that n=1025 and x=520

90% confidence interval for the Z = 1.645

a. population proportion p = x/n = 520/1025 = 0.507

b. margin of error = Z*sqrt(pq/n)
= 1.645*sqrt(0.507*0.493/1025)
= 0.0257

C. 90% confidece interval for the population proportion p = p +/- E

= 0.507 +/- 0.0257
= 0.4813 , 0.5327